# Optimizing a-DCF for Spoofing-Robust Speaker Verification

### v1

@@ III-C Baseline (S1)
TABLE II: Comparison of S1 and S2 in development set. Batch size is selected as 1024.
System min a-DCF
S2 0.1355
S3 0.1182
@@ III-D Proposed a-DCF metric optimization (S2, S3 and S4)
We next consider models S2 and S3, beginning with the hard-coded detection threshold of $\tau_{\mathrm{sasv}}=0.5$ in Eq. (5). Model selection is based on the lowest a-DCF value at a threshold of $0.5$ in the development set. The soft a-DCF only system (S2) attains minimum a-DCF of $0.1355$ and the combination of the soft a-DCF and BCE (S3) achieves minimum a-DCF of $0.1182$ as summarized in Table II.
TABLE III: Comparison of S1, S2, S3, and S4 in development and evaluation set. The results are given using 1024 as batch size.
Baseline (S1) S2 S3 S4
Dev 0.1234 0.1355 0.1182 0.1109
Eval 0.1445 0.2352 0.1398 0.1254
The results of S2 and S3 were based on hard-coded arbitrary detection threshold of $\tau_{\mathrm{sasv}}=0.5$. We now also optimize $\tau_{\mathrm{sasv}}$ using the complete procedure (Algorithm 1). Joint optimization of the DNN parameters and the threshold leads to noticeable improvements, as expected. Specifically, S4 achieved a minimum a-DCF of $0.1109$ on the development set and $0.1254$ on the evaluation set, outperforming S1, S2, and S3 as Table III indicates.
Figure 2: The plot illustrates the a-DCF values with respect to the threshold for the four different systems. The curves represent the a-DCF values for Baseline (S1), S2, S3, and S4 systems across various threshold values. Circular markers indicate the a-DCF values at a fixed threshold of $0.5$ for each system. Cross markers denote the minimum a-DCF values and their corresponding threshold points. The star marker represents the optimized threshold and the corresponding a-DCF value proposed for the S4 system.
Fig. 2 illustrates the a-DCF curves obtained for the four different systems (S1, S2, S3, and S4) as a function of the threshold. For systems S2 and S3, the a-DCF values at both the fixed threshold value of $0.5$ and their respective minimum points are depicted on the graph. For the S4 system, that includes threshold optimization, the a-DCF value corresponding to the optimized threshold is marked with a star. As can be seen from the graph (and supported by the results), the S4 model provides the lowest a-DCF value.
@@ III-E Impact of a-DCF Parameter
Figure 3: Target vs Nontarget and Target vs Spoof DET curves for three different a-DCF parameter settings. The SV-EER values for Setting1 (represented by the solid blue line), Setting1 (soft a-DCF + BCE) (represented by the dotted blue line), Setting2, and Setting3 are 13.97%, 8.44%, 7.52%, and 49.13%, respectively.
Our final analysis addresses the impact of the a-DCF parameters. To this end, we introduced three different set of a-DCF parameters as:
@@ IV Conclusion
In this study, we demonstrated that a spoofing-aware speaker verification system can be effectively optimized for the new a-DCF metric. By incorporating both the a-DCF and binary cross-entropy losses in the optimization process, we fine-tuned the network weights, which led to improved performance over the baseline system. Importantly, the usage of a novel threshold optimization technique further enhanced the system’s capabilities, resulting in a significant improvement in a-DCF values. These findings highlight the potential for adjusting ASV systems to achieve an optimal balance between user convenience and security against to evolving spoofing attacks.

### v2

@@ III-D Proposed a-DCF metric optimization (S2, S3 and S4)
We next consider models S2 and S3, beginning with the hard-coded detection threshold of $\tau_{\mathrm{sasv}}=0.5$ in the soft a-DCF loss equation. Model selection is based on the lowest a-DCF value at a threshold of $0.5$ in the development set. The soft a-DCF only system (S2) attains a minimum a-DCF of $0.1355$, with an SV-EER of $9.76\%$ and an SPF-EER of $0.14$ in the development set, while in the evaluation set, it reaches a minimum a-DCF of $0.2352$, an SV-EER of $17.28\%$ and an SPF-EER of $1.02\%$. In contrast, the combination of the soft a-DCF and BCE (S3) achieves a minimum a-DCF of $0.1182$, with an SV-EER of $8.29\%$ and an SPF-EER of $0.10\%$ in the development set, and in the evaluation set, it attains a minimum a-DCF of $0.1398$, an SV-EER of $9.59\%$ and an SPF-EER of $0.63\%$, indicating improved accuracy.
TABLE II: Comparison of S1, S2, S3, and S4 in development and evaluation set. The results are given using 1024 as batch size.
System a-DCF SV-EER SPF-EER
Dev Eval Dev Eval Dev Eval
Baseline (S1) 0.1234 0.1445 8.36 9.86 0.07 0.63
S2 0.1355 0.2352 9.76 17.28 0.14 1.02
S3 0.1182 0.1398 8.29 9.59 0.10 0.63
S4 0.1109 0.1254 7.75 8.44 0.08 0.61
The results of S2 and S3 were based on the hard-coded arbitrary detection threshold of $\tau_{\mathrm{sasv}}=0.5$. We now also optimize $\tau_{\mathrm{sasv}}$ using the complete procedure (Algorithm 1). Joint optimization of the DNN parameters and the threshold leads to noticeable improvements, as expected. Specifically, S4 achieved a minimum a-DCF of $0.1109$, with an SV-EER of $7.75\%$ and an SPF-EER of $0.08$ on the development set, and in the evaluation set, it reached a minimum a-DCF of $0.1254$, an SV-EER of $8.44\%$ and an SPF-EER of $0.61\%$, outperforming S1, S2, and S3, as Table II indicates.
Figure 2: a-DCF values:Baseline (S1), S2, S3, and S4 across different thresholds.
Fig. 2 presents the a-DCF curves for four systems (S1, S2, S3, and S4) across varying thresholds. For S2 and S3, both the a-DCF values at the fixed threshold of $0.5$ and the minimum points are indicated. The S4 system, which incorporates threshold optimization, has its optimized a-DCF value marked with a star. As seen in the graph, S4 yields the lowest a-DCF value.
@@ III-E Impact of a-DCF Parameter
Figure 3: Target vs. Nontarget and Target vs. Spoof DET curves for three a-DCF parameter settings. SV-EER values are 13.97% for Setting1 (solid blue), 8.44% for Setting1 (soft a-DCF + BCE) (dotted blue), 7.52% for Setting2, and 49.13% for Setting3.
Our final analysis addresses the impact of the a-DCF parameters. To this end, we introduced three different sets of a-DCF parameters as:
@@
III-F a-DCF Loss in Score Fusion
In this section, we demonstrated SASV performance using score fusion (SF) optimized with soft a-DCF. As reported in [21], nonlinear calibrated score fusion (NL-Cal-SF) performs better than both linear score sum (L-SF) and linear calibrated score sum (L-Cal-SF) fusion. Building on this, we applied soft a-DCF (aDCF-NL-Cal-SF) optimization with nonlinear calibrated score fusion. Performance metrics and trainable parameter counts are provided in Table III. The results indicate that our proposed SF approach improves min a-DCF from $0.0508$ to $0.0289$ compared to nonlinear calibrated SF. Additionally, we see that SF achieves better SASV performance compared with embedding fusion, despite using significantly fewer parameters ($4$ compared to $180736$).
TABLE III: Comparison of SASV systems with proposed score fusion.
SASV System #Trainable Params min a-DCF
Embedding Fusion 180736 0.1234
L-SF 0 0.5311
L-Cal-SF 4 0.0648
NL-Cal-SF 4 0.0508
aDCF-NL-Cal-SF 4 0.0289
@@ IV Conclusion
In this study, we demonstrated that an SASV system either embedding or score fusion can be effectively optimized for the new a-DCF metric. Incorporating both a-DCF and BCE losses in the optimization improved performance for embedding fusion over the baseline system. Importantly, the usage of a novel threshold optimization technique further enhanced the system’s capabilities, resulting in improvement in a-DCF values. These findings highlight the potential for adjusting ASV systems to achieve an optimal balance between user convenience and security against to evolving spoofing attacks.